OpenStudy (anonymous):

A baseball thrown straight up into the air has height s(t)=6+48t-16t^2 feet after t seconds.Its velocity at time t is given by v(t)=48-32t. At what time(s) will the ball be 38 ft above the ground?

OpenStudy (dumbcow):

plug in s(t) = 38 solve for t using quadratic formula

OpenStudy (anonymous):

can you show me how to

OpenStudy (dumbcow):

set it equal to 0 16t^2 -48t +32 = 0 divide by 16 t^2 -3t +2 = 0 (t-2)(t-1) = 0 t = 1,2 haha it factors so no quadratic formula needed